the perimeter of a rectangle is 200 feet. the width is two-thirds as long as length. what are the dimensions of the rectangle
Answer by rfer(10432)

P=2w+2L
200=2(2/3L)+2L
200=4/3L+2L
200=10/3L
L=200/(10/3)
L=200*3/10
L=600/10
L=60 ft
2/3L=40 ft